THE HOME DEPOT, INC. CONSOLIDATED BALANCE SHEETS February 3, 2019 January 28, 2018 in millions, except per share data AssetsTHE HOME DEPOT, INC. CONSOLIDATED STATEMENTS OF EARNINGS $ Fiscal 2018 108,203 $ 71,043 37,160 Fiscal 2017 100,904 $ 66,548 31. What is the Long-Term Liabilities to Current Liabilities?

2. What is the Asset Turnover?

3. What are the Return on Assets?

4. Even though net earnings increased, net equity decreased because:

a. Inventories shrank.

b. The company made major purchases of its corporate stock.

c. The company paid down its long-term debt.

d. income tax increased
Answer #1
Particulars 2019 2018
1) Long term liabilities to Current Liabilities =(26807+1867)/16716 =(24267+2174)/16194
= Long term Liab + Other Long term Liab/Current Liab 1.72 1.63
2) Asset Turnover =108203/44003 =100904/44529
= Turnover / Total Assets 2.46 2.27
3) Return on assets =11121/44003 =8630/44529
= Net earnings / Total Assets 25.27% 19.38%
Note: Total Assets can also by Average of Total Assets of two year.
4) Answer is b) The company made major purchases of its corporate stock. This can be seen from increase in Treasury stock.